OpenStudy (afloridagirl):

The first-, second-, and third-year enrollment values for a technical school are shown in the table below. Enrollment at a Technical School Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 2010 740 785 740 2011 690 710 781 2012 732 732 710 2013 781 755 800 Which of the following statements is true based on the data in the table?

OpenStudy (afloridagirl):

@agent0smith im sorry to tag you again but could help with this one...?

OpenStudy (agent0smith):

Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 2010 740 785 740 2011 690 710 781 2012 732 732 710 2013 781 755 800 Which of the following statements is true based on the data in the table? (where are the statements?)

OpenStudy (afloridagirl):

Which of the following statements is true based on the data in the table? The solution to f(x) = t(x) is x = 781. The solution to f(x) = t(x) is x = 2,011. The solution to s(x) = t(x) is x = 756. The solution to s(x) = t(x) is x = 2,009.

OpenStudy (afloridagirl):

@agent0smith

OpenStudy (afloridagirl):

i got it wrong, so i have a new question, can you help with it?

OpenStudy (agent0smith):

Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 ^^^s(x) and t(x) are equal in 2009

OpenStudy (agent0smith):

Sure

OpenStudy (afloridagirl):

im sorry

OpenStudy (afloridagirl):

The first-, second-, and third-year enrollment values for a technical school are shown in the table below. Enrollment at a Technical School Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 2010 740 785 740 2011 690 710 781 2012 732 732 710 2013 781 755 800 Which of the following statements is true based on the data in the table? The solution to f(x) = s(x) is x = 2,012. The solution to f(x) = s(x) is x = 732. The solution to s(x) = t(x) is x = 2,011. The solution to s(x) = t(x) is x = 710.

OpenStudy (afloridagirl):

i believe it is the second one

OpenStudy (agent0smith):

2012 732 732 710 f(x) = s(x) in 2012

OpenStudy (afloridagirl):

so..?

OpenStudy (afloridagirl):

never mind i was wrong

OpenStudy (agent0smith):

lol the one that has this f(x) = s(x) in 2012

OpenStudy (agent0smith):

The solution to f(x) = s(x) is x = 2,012.